# Relax Image-Specific Prompt Requirement in SAM: A Single Generic Prompt for Segmenting Camouflaged Objects

### 13 Dec 2023 v2

@@ Abstract
Camouflaged object detection (COD) approaches heavily rely on pixel-level annotated datasets. Weakly-supervised COD (WSCOD) approaches use sparse annotations like scribbles or points to reduce annotation effort, but this can lead to decreased accuracy. The Segment Anything Model (SAM) shows remarkable segmentation ability with sparse prompts like points. However, manual prompt is not always feasible, as it may not be accessible in real-world application. Additionally, it only provides localization information instead of semantic one, which can intrinsically cause ambiguity in interpreting the targets. In this work, we aim to eliminate the need for manual prompt. The key idea is to employ Cross-modal Chains of Thought Prompting (CCTP) to reason visual prompts using the semantic information given by a generic text prompt. To that end, we introduce a test-time adaptation per-instance mechanism called Generalizable SAM (GenSAM) to automatically enerate and optimize visual prompts the generic task prompt for WSCOD. In particular, CCTP maps a single generic text prompt onto image-specific consensus foreground and background heatmaps using vision-language models, acquiring reliable visual prompts. Moreover, to test-time adapt the visual prompts, we further propose Progressive Mask Generation (PMG) to iteratively reweight the input image, guiding the model to focus on the targets in a coarse-to-fine manner. Crucially, all network parameters are fixed, avoiding the need for additional training. Experiments demonstrate the superiority of GenSAM. Experiments on three benchmarks demonstrate that GenSAM outperforms point supervision approaches and achieves comparable results to scribble supervision ones, solely relying on general task descriptions as prompts.
@@
Figure 1: SAM encounters ambiguity and is sensitive to minor variations when using point and scribble prompts for segmentation. When using a generic task description as a generic prompt with CLIP Surgery+SAM enables the model to achieve some segmentation capability on obvious objects. However, it struggles to perform well in complex environments with camouflage-like patterns. In contrast, our proposed GenSAM can adaptively convert the generic task description into image-specific visual prompts, effectively enhancing the segmentation process by leveraging the unique characteristics of each input image.
@@ Introduction
But SAM has limitations in comprehending the content it segments. Manual prompts can only provide location information of the desired segmentation objects, but lack their precise semantic information, leading to potential ambiguity in their interpretation. e.g., In Fig.1(a), despite both prompts targeting the same object, minor changes in the point prompt’s position make SAM misinterpret the desired object, greatly altering the results. Moreover, compared to human perception, it also exhibits a strong bias in interpreting prompts (Chen et al. 2023a). In Fig.1(b), even with more information from scribble prompts than point prompts, SAM still misunderstands the target, leading to poorer performance. To relieve the ambiguity and bias, recent work expands manual prompt input options to include reference regions, videos and even audios (Zou et al. 2023). However, regardless of the prompting method, SAM still requires precise instance-specific manual prompts, which may not always be practical in real-world scenarios, and the question of eliminating this need remains unexplored.
In this work, we introduce a test-time adaptation mechanism called Generalizable SAM (GenSAM), a novel approach aiming to alleviate the strict demand for accurate, instance-specific manual prompts in the SAM framework for the WSCOD task. Given a simple text description, detailed semantic information about the desired object is reasoned based on both text and image information. Subsequently, this generates unambiguous visual prompt to guide the segmentation without human intervention. In order to provide semantic information of the target objects for SAM, we introduce Cross-modal Chains of Thought Prompting (CCTP) which automatically reason pixel-level visual prompts from various chains. BLIP2 (Zhu et al. 2023) and our devised CLIP (Radford et al. 2021) are employed for this propose. BLIP2 generates various keywords related to the target and their background using multiple chains of prompting (Wei et al. 2022), ambiguously in the input text prompts is eliminated. The spatial CLIP component introduces a novel self-attention mechanism, mapping keywords from different chains onto a consensus heatmap for generating consensus visual prompts, thus resolving the visual prompt ambiguity induced by a particular chain. To progressively adapt the visual prompts, we further propose Progressive Mask Generation (PMG), which uses an test-time prompt tuning approach to iteratively reweight the input image with the consensus heatmaps. This guides the model to focus on the targets in a coarse-to-fine manner. It encourages the model to concentrate on task-relevant regions, enhancing the performance.
@@
1). To eliminate the need for specific annotations tailored to each image in WSCOD, our GenSAM approach automatically generates personalized prompts for multiple unlabeled images using only a general task description.
2). To convert task descriptions into precise visual prompts, we introduce an Cross-modal Chains of Thought Prompting module. It uses a consensus mechanism and a novel self-attention to derive image-specific prompts for SAM. Additionally, Progressive Mask Generation module utilizes the consensus heatmap as a visual prompt, progressively enhancing the segmentation performance.
3). Extensive experiments on three benchmarks have demonstrated the effectiveness of our proposed GenSAM.
@@
Concealed Object Segmentation. The goal of camouflage detection is to identify objects that blend into complex backgrounds. Initially, some studies utilize low-level features such as texture, brightness, and color to distinguish the foreground from the background (Pike 2018; Hou and Li 2011; Sengottuvelan, Wahi, and Shanmugam 2008). Recently, several end-to-end approaches (Fan et al. 2020b, a) are proposed that achieve superior performance. However, most of these methods require fully annotated samples for training, which imposes a significant annotation burden. Weakly-Supervised Concealed Object Detection (WSCOD) aims to train a segmentation model using sparsely annotation like points and scribbles. Although WSCOD alleviates the reliance on pixel-level annotations, its performance is still limited by the quality and diversity of the training data. The lack of representative samples in a single dataset, as well as the restricted coverage of various scenes and objects, hinder the generalization ability of the model. For example, (He et al. 2023b) only requires scribble supervision, which achieves decent segmentation performance with lower annotation requirements per image. But its performance is limited by the quality of the annotations and lacks in generalization ability. Furthermore, to achieve satisfied performance across different datasets, current WSCOD approaches still requires separate training on different datasets, which limits their generalization ability.
@@
Segment Anything Model. Segment Anything Model (SAM) (Kirillov et al. 2023) is trained on the extensive SA-1B dataset, which comprises a vast collection of 11 million images and over 1 billion masks. This extensive dataset enables SAM to establish a robust foundation model for image segmentation with strong zero-shot generalization ability. While SAM is good at segmenting images, it struggles with segmenting camouflaged objects (Tang, Xiao, and Li 2023; Ji et al. 2023a, b). Moreover, its impressive ability requires the use of carefully crafted prompts to guide segmentation, which can be subjective and unclear. To address the challenges SAM encounters in camouflaged object detection, SAM-adaptor (Chen et al. 2023b) leverages a fully supervised dataset of camouflaged objects to train the encoder, yielding favorable results. However, this approach is hampered by its substantial demand for pixel-level annotated data. On the other hand, PLFMG (He et al. 2023b) enhances SAM’s performance in WSCOD task through the application of pseudo labeling and multi-scale feature grouping. Regrettably, this method remains contingent upon separate training for different datasets within the WSCOD task, indicative of a deficiency in robust generalization capabilities. In contrast, our proposed approach only mandates a generic task description, enabling us to perform effective segmentation of concealed objects in unsupervised images across diverse datasets within the WSCOD task through instance-level test-time adaptation.
@@
Test-time Adaptation aims to adapt the model to a test domain that exhibits a domain gap with the training data (Wang et al. 2020; Hu et al. 2020), in order to improve the performance on the test data (Niu et al. 2022). Currently, there are two main categories of test-time domain adaptation: backward-based adaptation and backward-free adaptation. The former often utilizes self-supervised learning methods to learn the data characteristics of the target domain through entropy minimization (Wang et al. 2020; Hu et al. 2019, 2022). The latter mostly achieves backward-free adaptation through batch normalization statistic adaptation. DUA (Mirza et al. 2022) employs a running average technique to update the statistics and achieve adaptation, while DIGA (Wang et al. 2023) utilizes distribution adaptation via batch normalization to effectively perform semantic segmentation under domain gaps. In our work, we employ instance-level test-time domain adaptation, which simply relies on a general task description for camouflage object segmentation. It allows accurate camouflage object segmentation across diverse datasets without sample-level supervision.
@@
Due to CLIP’s powerful open-vocabulary capability, it can handle various text descriptions. Therefore, we input the generated keywords ${fk}_{i}^{j}$ and $A_{j}^{back}$ into CLIP, aiming to leverage its cross-modal alignment capability to highlight the corresponding regions in the image related to the interested object. However, CLIP’s open-vocabulary capability also results in the generated heatmap containing numerous irrelevant information unrelated to the task.
@@
To further enhance the location accuracy of the heatmap and effectively explore the internal structures and semantic correlations within the image, we propose the k-k-v self-attention paralleled to the original k-q-v path. The element-wise multiplication of the keys vectors reduces interference from redundant features, enabling the self-attention mechanism to focus more on internal correlations within the input image. This results in a better representation of the image’s internal structure and patterns. Additionally, the “kkv” approach, using different vectors for values, preserves more information from the original inputs, enriching the context and enhancing the model’s expressive capabilities. For the $m$-th block of the visual encoder in CLIP, its input features are denoted as $S_{m}$, with $K=S_{m}\cdot W_{k}$, $V=S_{m}\cdot W_{v}$, and $Q=S_{m}\cdot W_{q}$, where $W_{k}\in\mathbb{R}^{d\times d_{k}}$, $W_{v}\in\mathbb{R}^{d\times d_{v}}$, and $W_{q}\in\mathbb{R}^{d\times d_{q}}$ are learnable parameter matrices.
Next, we split $K$, $V$, and $Q$ into $h_{0}$ individual heads, each with dimensions $d_{k}$, $d_{v}$, and $d_{q}$, respectively. Assuming we have $h_{0}$ heads, the dimensions of $K$, $V$, and $Q$ can be represented as follows: $K=\{k_{1},k_{2},...,k_{h_{0}}\}$, $Q=\{q_{1},q_{2},...,q_{h_{0}}\}$, and $V=\{v_{1},v_{2},...,v_{h_{0}}\}$. Here, $k_{n}\in\mathbb{R}^{d_{k}}$, $v_{n}\in\mathbb{R}^{d_{v}}$, and $q_{n}\in\mathbb{R}^{d_{q}}$ represent the transformation results of the $n$-th head. In our proposed k-k-v self-attention strategy, Q is replaced by K, and the expression for k-k-v self-attention is as follows:
@@
The output of the modified k-k-v self attention block $\hat{S}_{m+1}$ can be defined as follows:
$\hat{S}_{m+1}=\begin{cases}\text{None},\text{if }m<d\\ f_{\text{FFN}}({\text{attention}_{kkv}(S_{m})}+S_{m})\quad\text{if }m=d\\ f_{\text{FFN}}({\text{attention}_{kkv}({S_{m})}}+\hat{S}_{m})\quad\text{if }m>d\end{cases},$ (5)
$\hat{S}_{m}$ and $S_{m}$ represent the output of k-k-v self-attention and k-q-v self-attention, respectively, for the m-th layer block. Shallow layers in the spatial CLIP focus more on low-level features and less on higher level concepts (e.g. meanings of input keywords). The original k-q-v attention mechanism is used in these layers. Deep layers start to represent higher level semantics like people and animals. Thus, for layers at depth d, $S_{m}(d-1)$ is utilized as the input for the new module. Beyond depth d, $\hat{S}_{m}(d-1)$ and ${S}_{m}(d-1)$from the previous block are both used as input. Following Clip Surgery, the value of d is set to 7. These features are then accumulated and mixed using the fully connected layer $f_{\text{FFN}}$. Image $X$, processed through the original image path and the k-k-v image path we designed, yields image features $F_{I}(X)$ and $\hat{F}_{I}(X)$, respectively. The corresponding text feature of the keyword $A^{fore}_{j}$ and $A^{back}_{j}$ are $F_{fk}^{j}(X)$ and $F_{bk}^{j}(X)$.
@@
After obtaining the corresponding features of different keywords, our objective is to identify a consensus among these features to locate specific regions of interest related to the task. Specifically, the consensus derived from foreground keyword generation is utilized to pinpoint the precise location of the camouflaged objects, while the consensus derived from background keyword generation helps eliminate interference from the background in object localization.
For foreground keywords, the feature dimensions of the foreground keyword $F_{fk}^{j}(X)$ and the image feature $\hat{F_{I}}(X)$ are $N_{i}\times 1\times C$, where $N_{i}$ represents the size of the image tokens, 1 represents the size of the text token, and $C$ represents the number of channels. Consequently, the foreground similarity vector obtained for the $j$-th keyword, denoted as $F_{fs}^{j}(X)$, is defined as follows:
$F_{fs}^{j}(X)=\frac{F_{fk}^{j}(X)}{||F_{fk}^{j}(X)||_{2}}\odot\frac{\hat{F_{I}}(X)}{||\hat{F_{I}}(X)||_{2}},$ (6)
where $\odot$ is element-wise multiplication. L2 normalization is applied across the channel dimension. To derive a consensus among various image features corresponding to different foreground keywords, $F_{fs}(X)$ can be obtained as follows:
$F_{fs}(X)=\frac{\sum_{j=1}^{J}(F_{fs}^{j}(X))}{J},$ (7)
@@
Similarly, we can also obtain the corresponding background consensus $F_{bs}(X)$ as:
$F_{bs}(X)=\frac{\sum_{j=1}^{J}(F_{bs}^{j}(X))}{J},$ (8)
The resulting similarity heatmap $SI(X)$ is then obtained as follows:
$SI(X)=sum(F_{fs}(X)-F_{bs}(X)),$ (9)
where $SI\in\mathbb{R}^{N_{i}\times 1}$. $SI$ is then upsampled using bilinear interpolation. After upsampling $SI(X)$ to match the original size of image $X$, the resulting output can be regarded as the consensus heatmap $H$ corresponding to $X$.
@@
In the last step, we locate the desired objects by generating a consensus heatmap $H$. From this heatmap, we select high-confidence points to guide the segmentation process in SAM. When directly using consensus-based prompts for segmentation, the results may not be satisfactory in complex backgrounds. However, a single inference can provide a consensus heatmap that has good task-guiding properties. The brightness of different positions in the heatmap is often positively correlated with the degree of task relevance. Hence, we use the heatmap as a visual prompt to reweight the original image and guide the model during test time adaptation. The weighted image $X^{\prime}$ is as follows:
@@
where $w_{pic}$ is a hyper-parameter, which we have assigned a value of 0.3. The weighted image $X^{\prime}$ is then used as input image for next iteration. In this way, we develop a circularly test-time adaptation framework which involves multiple iterations of inference in a coarse-to-fine manner.
Moreover, in subsequent iterations, we use the previous iteration’s mask to guide the segmentation by drawing bounding boxes as a post-process. We select the box with the highest Intersection over Union (IoU) value with the mask as our choice. It optimizes the current iteration and improves the consistency of segmentation results. The mask obtained at the $i$-th iteration is defined as $M_{i}$. To eliminate the impact of ambiguity caused by inconsistent prompts in each iteration, the mask obtained in each iteration is averaged. Finally, the selected iteration $\mathrm{i}^{*}$ is determined by selecting the iteration’s result that closely resembles the average mask across all iterations as follows:
@@
Scale $M\downarrow$ $F_{\beta}\uparrow$ $E_{\phi}\uparrow$ $S_{\alpha}\uparrow$
0.125 0.110 0.496 0.750 0.689
0.25 0.082 0.596 0.806 0.741
0.5 0.067 0.681 0.838 0.775
@@
To test our method in different scenarios, we have chosen challenging camouflaged object detection (COD) datasets to evaluate our GenSAM under various settings.
@@
Datasets. Camouflaged object detection tasks aim to identify organisms attempting to camouflage themselves from complex backgrounds. In this study, we selected three representative datasets containing samples of camouflage objects: CHAMELEON (Skurowski et al. 2018), CAMO (Le et al. 2019) and COD10K (Fan et al. 2021a). CHAMELEON dataset comprises 76 images sourced from the Internet for testing purposes. CAMO dataset consists of 1,250 images, with 1,000 images allocated for training and 250 images for testing. COD10K dataset encompasses a total of 3,040 training samples and 2,026 testing samples.
Baseline. We have compared the current SOTA weakly supervised segmentation methods, namely SAM (Kirillov et al. 2023), WSSA (Zhang et al. 2020), SCWS (Yu et al. 2021), TEL (Zhang et al. 2020), and SCOD (He et al. 2023b). Three distinct levels of supervision are introduced, including scribble supervision and point supervision, alongside our proposed task-generic prompt settings. Following the previous weakly supervised segmentation setting (He et al. 2023a), in terms of scribble supervision, it involves providing foreground and background supervision by drawing the primary structure of objects and background areas. Point supervision refers to the provision of separate points as supervision for the foreground and background. In our newly proposed generic task prompt, we provide a unified prompt description “the camouflaged animal” for all images. The model is required to independently convert this unified description into specific supervision to guide the segmentation process based on the characteristics of each image. Regarding SAM, we follow the suggested setup by (He et al. 2023a) that involves comparing two variants of SAM: SAM-S and SAM-P. They finetune the mask decoder of SAM through scribble and point supervision training data respectively. Both variants employ partial cross-entropy loss for training. Note that all the comparative methods we have employed in our study are trained on camouflage segmentation datasets and tested on a separate test set. However, our approach differs significantly as it does not make use of a training set at all. Instead, we directly utilize the test set for test-time adaptation. Based on the approaches used in previous studies (Fan et al. 2021a, 2020a), we employ four commonly used metrics for evaluating the segmentation performance. These metrics include Mean Absolute Error ($M$), adaptive F-measure ($F_{\beta}$) (Margolin, Zelnik-Manor, and Tal 2014), mean E-measure ($E_{\phi}$) (Fan et al. 2021b), and structure measure ($S_{\alpha}$) (Fan et al. 2017). It is worth noting that a smaller value of $M$ or larger values of $F_{\beta}$, $E_{\phi}$, and $S_{\alpha}$ indicate better segmentation performance.

### 18 Dec 2023 v3

@@ Abstract
Camouflaged object detection (COD) approaches heavily rely on pixel-level annotated datasets. Weakly-supervised COD (WSCOD) approaches use sparse annotations like scribbles or points to reduce annotation efforts, but this can lead to decreased accuracy. The Segment Anything Model (SAM) shows remarkable segmentation ability with sparse prompts like points. However, manual prompt is not always feasible, as it may not be accessible in real-world application. Additionally, it only provides localization information instead of semantic one, which can intrinsically cause ambiguity in interpreting targets. In this work, we aim to eliminate the need for manual prompt. The key idea is to employ Cross-modal Chains of Thought Prompting (CCTP) to reason visual prompts using the semantic information given by a generic text prompt. To that end, we introduce a test-time instance-wise adaptation mechanism called Generalizable SAM (GenSAM) to automatically generate and optimize visual prompts from the generic task prompt for WSCOD. In particular, CCTP maps a single generic text prompt onto image-specific consensus foreground and background heatmaps using vision-language models, acquiring reliable visual prompts. Moreover, to test-time adapt the visual prompts, we further propose Progressive Mask Generation (PMG) to iteratively reweight the input image, guiding the model to focus on the targeted region in a coarse-to-fine manner. Crucially, all network parameters are fixed, avoiding the need for additional training. Experiments on three benchmarks demonstrate that GenSAM outperforms point supervision approaches and achieves comparable results to scribble supervision ones, solely relying on general task descriptions.
@@
Figure 1: In SAM, manual point and scribble prompts suffer from ambiguity in interpreting targets and is sensitive to minor spatial variations. Using a generic task description as a generic prompt with CLIP Surgery+SAM enables the model to achieve some segmentation capability on obvious objects. However, it struggles to perform well in complex environments with camouflage-like patterns. In contrast, our proposed GenSAM can adaptively convert a generic task description into image-specific visual prompts, effectively enhancing the segmentation process by leveraging the unique characteristics of each image.
@@ Introduction
However, SAM has limited comprehension on its segmented object. Manual prompts can only provide location information of desired segmentation objects, but lack in semantic information, leading to potential ambiguity. In Fig. 1(a), despite both prompts targeting the same object, minor changes in the point prompt’s position can make SAM misinterpret the desired object, greatly changing the results. Moreover, compared to human perception, it also exhibits a strong bias in interpreting prompts (Chen et al. 2023a). In Fig. 1(b), even with more information from scribble prompts than point prompts, SAM still misunderstands the target, leading to limited performance. To eliminate ambiguity and bias, recent work expands manual prompt input options to include reference regions, videos and even audios (Zou et al. 2023). However, regardless of the prompting method, SAM still requires instance-specific manual prompts, which may not always be practical in real-world scenarios, and the question of eliminating this need remains unexplored.
In this work, we introduce a test-time adaptation mechanism called Generalizable SAM (GenSAM), a novel approach to alleviating the demand for accurate, instance-specific manual prompts in the SAM framework for the WSCOD task. Given a simple text description, detailed semantic information about the desired object is reasoned based on both text and image information. Subsequently, this generates unambiguous visual prompt to guide the segmentation without human intervention. In order to provide semantic information of the target objects for SAM, we introduce Cross-modal Chains of Thought Prompting (CCTP) which automatically reason pixel-level visual prompts from various chains. BLIP2 (Zhu et al. 2023) and our devised CLIP (Radford et al. 2021) are employed for this propose. BLIP2 generates various keywords related to the target and their background using multiple chains of prompting (Wei et al. 2022), ambiguously in the input text prompts is eliminated. The spatial CLIP component introduces a novel self-attention mechanism, mapping keywords from different chains onto a consensus heatmap for generating consensus visual prompts, thus resolving the visual prompt ambiguity induced by a particular chain. To progressively adapt the visual prompts, we further propose Progressive Mask Generation (PMG), which uses an test-time prompt tuning approach to iteratively reweight the input image with the consensus heatmaps. This guides the model to focus on the targets in a coarse-to-fine manner. It encourages the model to concentrate on task-relevant regions, enhancing the performance.
@@
(1) To eliminate the need for specific annotations tailored to each image in WSCOD, our GenSAM approach automatically generates personalized prompts for multiple unlabeled images using only a general task description.
(2) To convert task descriptions into precise visual prompts, we introduce an Cross-modal Chains of Thought Prompting module. It uses a consensus mechanism and a novel self-attention to derive image-specific prompts for SAM. Additionally, Progressive Mask Generation module utilizes the consensus heatmap as a visual prompt, progressively enhancing the segmentation performance.
(3) Extensive experiments on three benchmarks have demonstrated the effectiveness of our proposed GenSAM.
@@
The goal of camouflage detection is to identify objects that blend into complex backgrounds. Initially, some studies utilize low-level features such as texture, brightness, and color to distinguish the foreground from the background (Pike 2018; Hou and Li 2011; Sengottuvelan, Wahi, and Shanmugam 2008). Recently, several end-to-end approaches (Fan et al. 2020b, a) are proposed that achieve superior performance. However, most of these methods require fully annotated samples for training, which imposes a significant annotation burden. Weakly-Supervised Concealed Object Detection (WSCOD) aims to train a segmentation model using sparsely annotation like points and scribbles. Although WSCOD alleviates the reliance on pixel-level annotations, its performance is still limited by the quality and diversity of the training data. The lack of representative samples in a single dataset, as well as the restricted coverage of various scenes and objects, hinder the generalization ability of the model. For example, (He et al. 2023b) only requires scribble supervision, which achieves decent segmentation performance with lower annotation requirements per image. But its performance is limited by the quality of the annotations and lacks in generalization ability. Furthermore, to achieve satisfied performance across different datasets, current WSCOD approaches still requires separate training on different datasets, which limits their generalization ability.
Segment Anything model
@@
Segment Anything Model (SAM) (Kirillov et al. 2023) is trained on the extensive SA-1B dataset, which comprises a vast collection of 11 million images and over 1 billion masks. This extensive dataset enables SAM to establish a robust foundation model for image segmentation with strong zero-shot generalization ability. While SAM is good at segmenting images, it struggles with segmenting camouflaged objects (Tang, Xiao, and Li 2023; Ji et al. 2023a, b). Moreover, its impressive ability requires the use of carefully crafted prompts to guide segmentation, which can be subjective and unclear. To address the challenges SAM encounters in camouflaged object detection, SAM-adaptor (Chen et al. 2023b) leverages a fully supervised dataset of camouflaged objects to train the encoder, yielding favorable results. However, this approach is hampered by its substantial demand for pixel-level annotated data. On the other hand, PLFMG (He et al. 2023b) enhances SAM’s performance in WSCOD task through the application of pseudo labeling and multi-scale feature grouping. Regrettably, this method remains contingent upon separate training for different datasets within the WSCOD task, indicative of a deficiency in robust generalization capabilities. In contrast, our proposed approach only mandates a generic task description, enabling us to perform effective segmentation of concealed objects in unsupervised images across diverse datasets within the WSCOD task through instance-level test-time adaptation.
Test-time Adaptation
@@
Test-time domain adaptation aims to adapt the model to a test domain that exhibits a domain gap with the training data (Wang et al. 2020; Hu et al. 2020), in order to improve the performance on the test data (Niu et al. 2022). Currently, there are two main categories of test-time domain adaptation: backward-based adaptation and backward-free adaptation. The former often utilizes self-supervised learning methods to learn the data characteristics of the target domain through entropy minimization (Wang et al. 2020; Hu et al. 2019, 2022). The latter mostly achieves backward-free adaptation through batch normalization statistic adaptation. DUA (Mirza et al. 2022) employs a running average technique to update the statistics and achieve adaptation, while DIGA (Wang et al. 2023) utilizes distribution adaptation via batch normalization to effectively perform semantic segmentation under domain gaps. In our work, we employ instance-level test-time domain adaptation, which simply relies on a general task description for camouflage object segmentation. It allows accurate camouflage object segmentation across diverse datasets without sample-level supervision.
@@
Due to CLIP’s powerful open-vocabulary capability, it can handle various text descriptions. Hence, we input the generated keywords ${A}_{j}^{fore}$ and $A_{j}^{back}$ into CLIP, aiming to leverage its cross-modal alignment capability to highlight the corresponding regions in the image related to the interested object. However, CLIP’s open-vocabulary capability also results in the generated heatmap containing numerous irrelevant information unrelated to the task.
@@
To further enhance the location accuracy of the heatmap and effectively explore the internal structures and semantic correlations within the image, we propose the k-k-v self-attention paralleled to the original k-q-v path. The element-wise multiplication of the keys vectors reduces interference from redundant features, enabling the self-attention mechanism to focus more on internal correlations within the input image. This results in a better representation of the image’s internal structure and patterns. Additionally, the “kkv” approach, using different vectors for values, preserves more information from the original inputs, enriching the context and enhancing the model’s expressive capabilities. For the $m$-th block of the visual encoder in CLIP, its input features are denoted as $S_{m}$, where $S_{m}\in\mathbb{R}^{L\times d}$, with $K=S_{m}\cdot W_{k}$, $V=S_{m}\cdot W_{v}$, and $Q=S_{m}\cdot W_{q}$. Here, $W_{k}\in\mathbb{R}^{d\times d_{k}}$, $W_{v}\in\mathbb{R}^{d\times d_{v}}$, and $W_{q}\in\mathbb{R}^{d\times d_{q}}$ are learnable parameter matrices. Next, we split $K$, $V$, and $Q$ into $h_{0}$ individual heads, each with dimensions $d_{k}$, $d_{v}$, and $d_{q}$, respectively. Assuming we have $h_{0}$ heads, the dimensions of $K$, $V$, and $Q$ can be represented as follows: $K=\{k_{1},k_{2},...,k_{h_{0}}\}$, $Q=\{q_{1},q_{2},...,q_{h_{0}}\}$, and $V=\{v_{1},v_{2},...,v_{h_{0}}\}$. Here, $k_{n}\in\mathbb{R}^{L\times d_{k}}$, $v_{n}\in\mathbb{R}^{L\times d_{v}}$, and $q_{n}\in\mathbb{R}^{L\times d_{q}}$ represent the transformation results of the $n$-th head. In our proposed k-k-v self-attention strategy, Q is replaced by K, and the expression for k-k-v self-attention is as follows:
@@
where $scale$ is the scaling factor. The output of the K-Q-V block ${S}_{m+1}$ and modified K-K-V self attention block $\hat{S}_{m+1}$ are defined as follows:
$\begin{split}&S_{m+1}=f_{\text{FFN}}(\text{attention}_{kqv}(S_{m})+S_{m}),\\ &\hat{S}_{m+1}=\begin{cases}\text{None},\text{if }m<\delta\\ f_{\text{FFN}}({\text{attention}_{kkv}(S_{m})}+S_{m})\quad\text{if }m=\delta\\ f_{\text{FFN}}({\text{attention}_{kkv}({S_{m})}}+\hat{S}_{m})\quad\text{if }m>\delta\end{cases},\end{split}$ (5)
where $\hat{S}_{m}$ and $S_{m}$ represent the output of k-k-v self-attention and k-q-v self-attention, respectively, for the m-th layer block. Shallow layers in the spatial CLIP focus more on low-level features and less on higher level concepts (e.g. meanings of input keywords). The original k-q-v attention mechanism is used in these layers. Deep layers start to represent higher level semantics like people and animals. Thus, for layers at depth $\delta$, $S_{m-1}$ is utilized as the input for the new module. Beyond depth $\delta$, $\hat{S}_{m-1}$ and ${S}_{m-1}$from the previous block are both used as input. Following Clip Surgery (Li et al. 2023), the value of $\delta$ is set to 7. These features are then accumulated and mixed using the fully connected layer $f_{\text{FFN}}$. Image $X$, processed through the original image path and the k-k-v image path, yields image features $F_{I}$ and $\hat{F}_{I}$, respectively. The corresponding text feature of the keyword $A^{fore}_{j}$ and $A^{back}_{j}$ are $F^{fore}_{j}$ and $F^{back}_{j}$.
@@
After obtaining the corresponding features of different keywords, our objective is to identify a consensus among these features to locate specific regions of interest related to the task. Specifically, the consensus derived from foreground keyword generation is utilized to pinpoint the precise location of the camouflaged objects, while the consensus derived from background keyword generation helps eliminate interference from the background in object localization. For foreground keywords, the feature dimensions of the foreground keyword $F^{fore}_{j}$ and the image feature $\hat{F_{I}}$ are $N_{i}\times 1\times C$, where $N_{i}$ represents the size of the image tokens, 1 represents the size of the text token, and $C$ represents the number of channels. Consequently, the foreground similarity vector $SI_{fore}^{j}$ obtained for the $j$-th keyword, is defined as:
$SI_{fore}^{j}=\frac{F^{fore}_{j}}{||F^{fore}_{j}||_{2}}\odot\frac{\hat{F_{I}}}{||\hat{F_{I}}||_{2}},$ (6)
where $\odot$ is element-wise multiplication. L2 normalization is applied across the channel dimension. To derive a consensus among various image features corresponding to different foreground keywords, $SI_{fore}$ can be obtained as follows:
$SI_{fore}=\frac{\sum_{j=1}^{J}(SI_{fore}^{j})}{J},$ (7)
@@
We also obtain the corresponding background consensus $SI_{back}$ in a similar way. Then the resulting similarity heatmap $SI$ is:
$SI=SI_{fore}-SI_{back},$ (8)
where $SI\in\mathbb{R}^{N_{i}\times 1}$. $SI$ is then upsampled using bilinear interpolation. After upsampling $SI$ to match the original size of image $X$, the resulting output can be regarded as the consensus heatmap $H$ corresponding to $X$. We further sample highly-activated pixels on $H$ as positive point prompts and the same number of the most unactivated pixels as negative point prompts to guide the segmentation process in SAM.
@@
However, a single inference may not provide satisfactory segmentation result. For image with complicated background, some backgound objects can also be highly activated in the heatmap, causing some noises for inference the point prompts. In order to get more roubust prompt, we use the heatmap as a visual prompt to reweight the original image and guide the model during test time adaptation. The weighted image $X^{\prime}$ is as follows:
@@
where $w_{pic}=0.3$ is a hyper-parameter. The weighted image $X^{\prime}$ is then used as input image for next iteration. In this way, we develop a circularly test-time adaptation framework which involves multiple iterations of inference in a coarse-to-fine manner.
Moreover, in subsequent iterations, we use the previous iteration’s mask to guide the segmentation by drawing bounding boxes as a post-process. We select the box with the highest Intersection over Union (IoU) value with the mask as our choice. It optimizes the current iteration and improves the consistency of segmentation results. The mask obtained at the $i$-th iteration is defined as $M_{i}$, where $i\in\{1,..,\mathbf{Iter}\}$, $\mathbf{Iter}$ is set as 6. To eliminate the impact of ambiguity caused by inconsistent prompts in each iteration, the mask obtained in each iteration is averaged. Finally, the selected iteration $\mathrm{i}^{*}$ is determined by selecting the iteration’s result that closely resembles the average mask across all iterations as follows:
@@
Factor $M\downarrow$ $F_{\beta}\uparrow$ $E_{\phi}\uparrow$ $S_{\alpha}\uparrow$
8 0.110 0.496 0.750 0.689
4 0.082 0.596 0.806 0.741
2 0.067 0.681 0.838 0.775
@@
To evaluate GenSAM in different scenarios, we choose challenging camouflaged object detection (COD) datasets to evaluate our GenSAM under various settings.
@@
Datasets. Camouflaged object detection tasks aim to identify organisms attempting to camouflage themselves from complex backgrounds. In this study, we select three representative datasets containing samples of camouflage objects: CHAMELEON (Skurowski et al. 2018), CAMO (Le et al. 2019) and COD10K (Fan et al. 2021a). CHAMELEON dataset comprises 76 images sourced from the Internet for testing purposes. CAMO dataset consists of 1,250 images, with 1,000 images allocated for training and 250 images for testing. COD10K dataset encompasses a total of 3,040 training samples and 2,026 testing samples.
Baseline. We compare current SOTA weakly supervised segmentation methods, namely SAM (Kirillov et al. 2023), WSSA (Zhang et al. 2020), SCWS (Yu et al. 2021), TEL (Zhang et al. 2020), and SCOD (He et al. 2023b). Three distinct levels of supervision are introduced, including scribble supervision and point supervision, alongside our proposed task-generic prompt settings. Following the previous weakly supervised segmentation setting (He et al. 2023a), in terms of scribble supervision, it involves providing foreground and background supervision by drawing the primary structure of objects and background areas. Point supervision refers to the provision of separate points as supervision for the foreground and background. In our newly proposed generic task prompt, we provide a unified prompt description “the camouflaged animal” for all images. The model is required to independently convert this unified description into specific supervision to guide the segmentation process based on the characteristics of each image. Regarding SAM, we follow the suggested setup by (He et al. 2023a) that involves comparing two variants of SAM: SAM-S and SAM-P. They finetune the mask decoder of SAM through scribble and point supervision training data respectively. Both variants employ partial cross-entropy loss for training. Note that all the comparative methods we employ in our study are trained on camouflage segmentation datasets and tested on a separate test set. However, GenSAM does not require training data at all, while directly utilize the test set for test-time adaptation. Based on the approaches used in previous studies (Fan et al. 2021a, 2020a), we employ four commonly used metrics for evaluation, include Mean Absolute Error ($M$), adaptive F-measure ($F_{\beta}$) (Margolin, Zelnik-Manor, and Tal 2014), mean E-measure ($E_{\phi}$) (Fan et al. 2021b), and structure measure ($S_{\alpha}$) (Fan et al. 2017). It is worth noting that a smaller value of $M$ or larger values of $F_{\beta}$, $E_{\phi}$, and $S_{\alpha}$ indicate better segmentation performance.